Find a basis for R(A)^⊥, where R(A) denotes the row space of the matrix
A where A is a 3x4 matrix (1 0 4 0 \ 0 1 2 -1 \ 1 -1 2 1)
1
Expert's answer
2020-03-31T05:29:03-0400

"R(A)^{\perp}=" a basis of the orthogonal complement of "R(A)."

Let "R(A)=\{ v_1,v_2,v_3 \}," where

"v_1=(1,0,4,0),v_2=(0,1,2,-1) \ \text{and} \ v_3=(1,-1,2,1)".

Since "v_2+v_3-v_1=0," the vectors "v_1, v_2, v_3" are linearly dependent. Thus, "\{v_1,v_2\}" is the maximal subset of linearly independent vectors in the row space.

If "w_i \in R(A)^{\perp}", "w_i=(a_i,b_i,c_i,d_i)," "v_1=(1,0,4,0) \in R(A)," "v_2=(0,1,2,-1) \in R(A)," then "w_i\cdot v_1=0,w_i \cdot v_2=0," hence "a_i+4c_i=0, b_i+2c_i-d_i=0."

If "c_1=1, d_1=0," then "a_1=-4c_1=-4," "b_1=-2c_1+d_1=-2," hence "w_1=(a_1,b_1,c_1,d_1)=(-4,-2,1,0)."

If "c_2=0,d_2=1," then "a_2=-4c_2=0," "b_2=-2c_2+d_2=1," hence "w_2=(a_2,b_2,c_2,d_2)=(0,1,0,1)."

Thus, "R(A)^{\perp}=\{w_1,w_2\}=\{ (-4,-2,1,0),(0,1,0,1)\}."